OpenStudy (anonymous):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem. Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he mad

OpenStudy (anonymous):

it is quite simple ... i guess

OpenStudy (anonymous):

awnser it then lol?@croslali

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

see what is the profit they make on 1 flower ?? it is 40 percent of the cost of a single flower right?

OpenStudy (anonymous):

ie 40/100*0.5

OpenStudy (anonymous):

understood

OpenStudy (anonymous):

write down whjat i would put down as my awnser

OpenStudy (anonymous):

hence for getting a profit of 500 dollars , let the number of flowers be sold be x then number of flowers by equation is (40/100)*(0.5)*x=500

OpenStudy (anonymous):

what wouldthe awnser go like on my assignment howwouldit be written

OpenStudy (anonymous):

that gives x=2500 flowers ... Max can identify if his equation is wrong if when he puts the value of 2500 in the equation above he should get 500 on the lhs if ot his answer is wrong

OpenStudy (anonymous):

okay write down everything u did soi can copy and paste it into my awnser

OpenStudy (anonymous):

okay this is the answer in whole profit of one flower=0.5*40/100 profit of x number of flowers which give profit of 500 dollars =0.5*(40/100)*x therefore the equation is that gives us the profit is (40/100)*(0.5)*x=500 that gives us , x=2500 flowers Max can identify if his equation is wrong if when he puts the value of 2500 in the equation above he should get 500 on the lhs if not his equation is wrong..

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

awnser the other ones i posted

OpenStudy (anonymous):

Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 8 = 7x−1 6 24 1 ( 3x+6 8 ) = 24 1 ( 7x−1 6 @croslali ) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x @croslali

OpenStudy (anonymous):

huh?!?!??!

OpenStudy (anonymous):

i posted it wrong figure out the one above ur name@croslali